# Fast algorithms to improve fair information access in networks

### v1

@@ III.2 Spreadability
Briefly, using ProbEst, given transmission probability $\alpha$, number of simulation rounds $R$, and a seed set $S$, we estimate $\pi_{i}$ for every $i\in V$ using $R$ independent cascades originating from $S$ Fish_2019 . The worst-case time complexity of ProbEst is $\mathcal{O}(R(|S|+2m))$ Fish_2019 , where $m$ is the number of edges in $G$. In practice, the computational cost of ProbEst increases both as the seed set grows and as $\alpha$ increases, because both changes tend to increase the size of the induced information cascades. Past work used $R=1000$ as a balance between statistical accuracy for $\pi_{i}$ and computational cost, which we incorporate in our work as well Fish_2019 . The output of ProbEst is used here both as part of some evaluated algorithms and to evaluate their performance.
@@
Prior work evaluated algorithm performance using transmission probabilities $\alpha\in\{0.3,0.4,0.5\}$ Fish_2019 . However, the resulting information cascades, and hence the associated access probabilities, are not a simple function of $\alpha$; they instead also depend on the network’s structure. For example, denser, more connected networks contain many more paths by which information can spread than do sparse networks. Hence the same $\alpha$ will tend to produce larger cascades on the former, and smaller cascades on the latter. To control for these structure-induced differences in information cascades, we introduce the concept of spreadability, which jointly accounts for the impact of network structure and transmission rate $\alpha$ on the sizes of information cascades.
@@
Figure 4: Minimum access probability $\pi_{\min}$ vs. seed set size $k$, with a best fit line (Myopic $\hat{\beta}=0.039$), averaged over 20 runs, evaluated on a large economic network ($n=2113$ nodes, $m=57927$ edges), with $\alpha=0.4$ and a budget of $k=10$ seeds, plus one random initial seed.
@@
We can construct a fine-grained spreadability function that relates a particular choice of $\alpha$ to the fraction of a network activated under the independent cascade model from a uniformly random initial seed. For a given $\alpha$, the spreadability $f(\alpha)$ on a particular network is computed by averaging over $R$ trials the fraction of nodes that are activated from a uniformly random initial seed. This calculation produces a monotonically increasing curve, as seen in Figure 3.
We find that computing spreadability for each $\alpha\in\{0.01,0.02,...0.99\}$ provides ample resolution to choose $\alpha$ close to target spreadabilities of $0.2,0.5,0.8$ (which we refer to as “low,” “medium,” and “high” spreadabilities respectively). We set $R=1000$, as the spreadability curve tends to stabilize near this value and we get diminishing returns for larger $R$.
III.3 New Metric for Algorithm Evaluation
We propose a metric for evaluating the relative performance of algorithms, $\beta$, which corresponds to the average slope of the lines of best fit for ProbEst-based evaluations after selecting the first $10$ seeds, as seen in Figure 4. This new metric is useful because it shows the marginal improvement in $\pi_{\rm min}$ that we can expect when asking an algorithm to draw a new seed, which allows for a direct comparison of the performance of the algorithms. We use this metric to evaluate the performance of our algorithms on the network corpus.
IV Algorithms for Fair Information Access
IV.1 Algorithms from Prior Work
@@
Four previously introduced heuristics for choosing seed nodes to maximize $\pi_{\rm min}$ are Greedy, Myopic, Naive Myopic and Gonzalez, along with Random as a baseline for comparison Fish_2019 . Random selects $k$ seeds by uniformly sampling nodes from the network. Given a partial seed set, Gonzalez selects as the next seed the node that is the furthest (in the shortest-path metric) from all nodes in the current set Fish_2019 . The Greedy algorithm iteratively selects a new seed by choosing the node with the highest marginal gain relative to the current seed set according to ProbEst. Due to its immensely high computational cost, Greedy is not practical for most networks Fish_2019 , and is not used in our study.
@@
In contrast, Myopic uses ProbEst with the current seed set, and selects as the next seed the node with the lowest $\pi_{i}$. Naive Myopic is similar to Myopic, but only runs ProbEst once, at initialization, and then selects as the seed set the nodes with the lowest $\pi_{i}$ values. In past work, on a small set of networks, Myopic was found to perform best. However, because Myopic depends on ProbEst, it is computationally expensive, which limits its applicability to large networks. In practice, Myopic is always the second slowest algorithm, after Greedy Fish_2019 .
Figure 5: Mean performance of the intervention algorithms on each domain in the corpus under medium spreadability. Each algorithm’s performance is averaged over a given domain, 20 runs per network.
@@
Myopic, Naive Myopic, and Gonzalez all start with an initial seed. Past work chose this initial seed to be the highest degree node. Our initial experiments indicate that this choice confers a substantial advantage to these methods (Fig. S1), and that some of the previous positive results are thus attributable to this initial seed choice rather than to the algorithms’ subsequent choices. To mitigate this bias we instead initialize all heuristics with a seed set composed of a single uniformly randomly selected node. Moreover, for each evaluation round on a particular network, we initialize all algorithms to use the same random seed, which controls for the effects of different initial seeds. This initial seed choice is not counted against the budget $k$ (Fig. 4).
IV.2 New Fast Algorithms
In addition to the four heuristics from prior work, here we introduce 10 new heuristics to maximize the minimum $\pi_{i}$ on a budget. These heuristics are designed to be computationally lightweight, scaling to far larger networks, while also matching or exceeding the performance of Myopic on the corpus. The new algorithms can be grouped into three families: BFS-based, PPR-based and Topology-based.
@@ IV.2 New Fast Algorithms
While this approach does not exactly measure $\pi_{i}$, it is much faster than ProbEst, taking $\mathcal{O}(n\cdot\langle k\rangle)$ per iteration, because for most networks, including those in our corpus, the mean degree $\langle k\rangle\ll 1000$ (Fig. 2; here, $k$ denotes the degree of a node). The key design principle of this algorithm is to capture complex network structures that Gonzalez could not account for. In Figure 4, we find that Myopic BFS almost matches Myopic’s performance in situations where Gonzalez lags behind.
@@
In the PPR-based heuristics, Myopic PPR and Naive Myopic PPR, we use Personalized PageRank (PPR) to estimate $\pi_{i}$ instead of ProbEst or BFS. Personalized PageRank (implemented with networkx SciPyProceedings_11 ) performs a random walk that probabilistically restarts from nodes in the seed set PPR1998 . The ranking produced by PPR is not a direct estimate of $\pi_{i}$; we treat the PPR values as being correlated, such that a lower PPR score is a proxy for a lower $\pi_{i}$ value.
Topology-based: LeastCentral and MinDegree variants
@@
The topology-based heuristics are based on the intuition that nodes with low values of $\pi_{i}$ have distinctive structural positions or patterns of connectivity. LeastCentral selects the non-seed node $i$ with the lowest closeness centrality $c_{i}$ as the next seed. Similarly, LeastCentral_n selects the lowest centrality node $i$’s highest degree neighbor as the next seed. Here, the closeness centrality of a node $i$ is the inverse of the average shortest path length from $i$ to all other reachable nodes in the same connected component FREEMAN1978215 ; newman18 . Lower closeness centrality implies the node is less reachable by the rest of the network, and therefore is expected to have lower $\pi_{i}$.
@@
The four remaining topology-based heuristics exploit a network’s degree structure to make decisions, based on the observation that in practice Myopic tends to select nodes with low degree as seeds. The first two heuristics take a non-seed node with the lowest degree, breaking ties by choosing the node with the lowest harmonic centrality boldi2013axioms ; newman18 among same-degree nodes. In the case of MinDegree_hc, it chooses that node itself as the next seed, while MinDegree_hcn chooses the highest-degree neighbor of that node. The two remaining variations replace harmonic centrality in the aforementioned logic with the neighbor degree, i.e. breaking ties by choosing the node with the highest neighbor degree (sum of degrees across neighbors). These heuristics are called MinDegree_nd and MinDegree_ndn, respectively.
@@
Figure 6: Performance of intervention algorithms on the network corpus, relative to Myopic and sorted in ascending order by network size within each domain. Under medium spreadability, $24\%$ of networks have no algorithm better than or within $80\%$ of Myopic’s performance. “Equivalent” is defined as within one standard error of $\beta$ for Myopic; typically about 0.001.
V Experimental Results
V.1 Performance of Algorithms on the Corpus
@@
We evaluate and compare the performance of 14 algorithms in total (10 new algorithms and 4 from prior work), applied to all $174$ networks in the corpus. We are interested in algorithms that operate well on a tight budget and so let $k=10$ seed nodes. For each of the three spreadability levels (low, medium, high), we produce a $10\times 14\times 174$ matrix, where each entry is the $\beta$ performance of an algorithm after adding $k_{n}\in\{1,\dots,10\}$ seeds in a network, averaged over $20$ runs. We focus on the medium spreadability here, and include results for low and high spreadability in the appendix. Figure 5 displays the mean performance of each algorithm on networks in a given domain under medium spreadability (see Fig. S2 for low and high spreadability results). Across domains, Myopic produces the best average performance.
@@
To further examine the results, we sort within each domain by network size in ascending order, and then score algorithms as better than, “equivalent” to (within one std. err.), within 80% of, or worse than Myopic (see Fig. 6; low and high spreadability results in Fig. S3). In this way, we can assess whether the better average performance of Myopic applies to individual networks compared to other algorithms. This experiment reveals that Myopic is not universally the best algorithm for all networks in any spreadability setting; on many individual networks other algorithms perform equivalently or better. In the medium spreadability setting, for only $24\%$ of networks is there no algorithm that performs at least $80\%$ as well as Myopic (Fig. 6).
@@
This variability in performance across networks suggests that network structure plays a critical role in governing the relative performance of different algorithms. Figure 7 plots the instances for which each algorithm was the best performing algorithm on a network against that network’s mean degree $\langle k\rangle$. Myopic tends to perform better on networks with lower average degree, although it does also perform well for many networks with larger mean degree. In practice, we find that the final seed set produced by Myopic is often composed primarily of low-degree nodes located in a network’s periphery far from one another, and these nodes may be too far removed from other disadvantaged nodes to meaningfully improve their $\pi_{i}$ values.
@@
A second takeaway is that a few specific algorithms tend to perform better than Myopic in certain settings (Fig. 7), specifically MinDegree_hcn and Gonzalez. The performance of MinDegree_hcn in particular tends to improve over Myopic with increasing average degree, while Gonzalez does best in networks with very low mean degrees. Furthermore, we note that in Figure S2A, on average MinDegree_hcn outperforms Myopic in the economic domain, and from Table S1, we see that the economic domain has the highest mean degrees. The MinDegree_hcn algorithm selects as seeds the highest-degree neighbors of low-degree nodes. As result, new information cascades seeded at these nodes will tend to spread quickly to a number of disadvantaged nodes in the network.
@@
Figure 7: Best-performing algorithm vs. mean degree of the network (medium spreadability), for all networks. Counts on the right show total circles per line, i.e., the number of times an algorithm was the best over the whole corpus.
V.2 Algorithm Runtime
@@
All of the new algorithms are substantially faster than Myopic and Naive Myopic (Fig. 8). Because they only require sorting two lists, MinDegree_ndn and MinDegree_nd are the most efficient, improving running times over ProbEst-based algorithms by a factor of 1000-10000x, depending on the size of the network (Fig. 8). BFS-based algorithms are marginally slower than these fastest algorithms, and algorithms that use an All-Pairs-Shortest-Paths (APSP) calculation fall between the ProbEst and BFS algorithm groups. As we would expect from the asymptotics, BFS algorithms tend to scale more slowly (in terms of runtime) with network size than do ProbEst algorithms, while APSP algorithms scale more quickly (Fig. 8).
The low upfront cost of APSP-based algorithms makes them far faster than ProbEst-based algorithms in many practical settings, being 10-100x faster on networks with less than $n=10^{6}$. However, the asymptotic cost of APSP grows super-linearly, implying that for sufficiently large networks, ProbEst will be faster. For our corpus, we estimate the crossover point when Myopic becomes more efficient than MinDegree_hc to occur between $1,262,000$ and $1,515,000$ nodes (95% CI, from 1000 bootstraps). However, approximation algorithms for APSP could potentially extend their practical efficiency much further.
Figure 8: Runtime of old and new algorithms across the network corpus with $k=10$ seeds averaged over 10 runs, showing a substantial advantage in running time for the new algorithms. Note: algorithms with * require ProbEst.
V.3 A Fast Meta-Learning Algorithm

### v2

@@ 2.2. Spreadability
Briefly, using ProbEst, given transmission probability $\alpha$, number of simulation rounds $R$, and a seed set $S$, we estimate $\pi_{i}$ for every $i\in V$ using $R$ independent cascades originating from $S$ (Fish_2019, ). The worst-case time complexity of ProbEst is $\mathcal{O}(R(|S|+2m))$ (Fish_2019, ), where $m$ is the number of edges in $G$. In practice, the computational cost of ProbEst increases both as the seed set grows and as $\alpha$ increases, because both changes tend to increase the size of the induced information cascades. Past work used $R=1000$ as a balance between statistical accuracy for $\pi_{i}$ and computational cost, and we follow this precedent (Fish_2019, ). Here, ProbEst is used as a subroutine in some algorithms, as well as to evaluate the performance of algorithms by estimating the achieved minimum activation probability of a computed seed set.
Selecting $\alpha$ values.
@@
Prior work evaluated algorithm performance using transmission probabilities $\alpha\in\{0.3,0.4,0.5\}$ (Fish_2019, ). However, the resulting information cascades, and hence the associated access probabilities, are not a simple function of $\alpha$; they instead also depend on the network’s structure. For example, denser, more connected networks contain many more paths by which information can spread than do sparse networks. Hence the same $\alpha$ will tend to produce larger cascades on the former, and smaller cascades on the latter. To control for these structure-induced differences in information cascades, we introduce the concept of spreadability, which jointly accounts for the impact of network structure and transmission rate $\alpha$ on the sizes of information cascades.
@@
Figure 4. Minimum access probability $\pi_{\min}$ vs. seed set size $k$, with a best fit line (Myopic $\hat{\beta}=0.039$), averaged over 20 runs, evaluated on a large economic network ($n=2113$ nodes, $m=57927$ edges), with $\alpha=0.4$ and a budget of $k=10$ seeds, plus one random initial seed.
@@
We can construct a fine-grained spreadability function that relates a particular choice of $\alpha$ to the fraction of a network activated under the independent cascade model from a uniformly random initial seed. For a given $\alpha$, the spreadability $f(\alpha)$ on a particular network the fraction of nodes that are activated from a uniformly random initial seed, averaged over $R$ trials. This calculation produces a monotonically increasing curve, as seen in Figure 3. We find that computing spreadability for each $\alpha\in\{0.01,0.02,...0.99\}$ provides ample resolution to choose $\alpha$ close to target spreadabilities of $0.2,0.5,0.8$ (which we refer to as “low,” “medium,” and “high” spreadabilities respectively). We set $R=1000$, as the spreadability curve tends to stabilize near this value and we get diminishing returns for larger $R$.
2.3. New Metric for Algorithm Evaluation
As discussed in Section 1.2, it is not straightforward to compare the performance of algorithms, even with a fixed network and $\alpha$ value. There are several reasons for this. First, we do not want to make strong assumptions about the “most useful” value for the parameter $k$ (the number of seeds to be added), and it is possible that algorithm $\mathcal{A}$ obtains a better objective score (the minimum activation probability $\pi_{\text{min}}$) than algorithm $\mathcal{A}^{\prime}$ for one value of $k$, but not for another. Moreover, many algorithms (in particular those using ProbEst) have some randomness. Finally, even our analysis of the quality of a solution is inexact; recall that exact computation of the $\pi_{\text{min}}$ achieved by a seed set $S$ is #P-hard (chen2010scalable, ). Thus, our ability to compare two solutions is limited by the precision of ProbEst.
We overcome these challenges by introducing a metric $\beta$, which intuitively measures the marginal gain in the problem objective (the minimum activation probability $\pi_{\text{min}}$) which we can expect when asking an algorithm to add one additional seed. Formally, for a given algorithm $\mathcal{A}$, network $G$, independent cascade parameter $\alpha$, and budget $k$, we run $\mathcal{A}$ on $(G,\alpha,k)$ and record a value $\pi_{\text{min}}(k)$ indicating the achieved objective score. We repeat for each $k\in\{1,2,\ldots,10\}$, computing the points $(k,\pi_{\text{min}}(k))\in\mathbb{N}\times[0,1]$. We then record the slope $\beta_{1}$ of the line of best fit for these points. We repeat the entire process multiple (generally $20$) times, producing slopes $\beta_{1},\beta_{2},\ldots$. The metric $\beta$ is the mean of these values, i.e., the average slope of the line of best fit; see Figure 4. Henceforth, when comparing the performance of algorithms, we do so via the metric $\beta$.
3. Algorithms for Fair Information Access
3.1. Algorithms from Prior Work
@@
Four previously introduced heuristics for choosing seed nodes to maximize $\pi_{\rm min}$ are Greedy, Myopic, Naive Myopic and Gonzalez, along with Random as a baseline for comparison (Fish_2019, ). Random selects $k$ seeds by uniformly sampling nodes from the network. Given a partial seed set, Gonzalez selects as the next seed the node that is the furthest (in the shortest-path metric) from all nodes in the current set (Fish_2019, ). The Greedy algorithm iteratively selects a new seed by choosing the node with the highest marginal gain relative to the current seed set according to ProbEst. Due to its immensely high computational cost, Greedy is not practical for most networks (Fish_2019, ), and is not used in our study.
@@
In contrast, Myopic uses ProbEst with the current seed set, and selects as the next seed the node with the lowest $\pi_{i}$. Naive Myopic is similar to Myopic, but only runs ProbEst once, at initialization, and then selects as the seed set the $k$ nodes with the lowest $\pi_{i}$ values. In past work, on a small set of networks, Myopic was found to perform best. However, because Myopic depends on ProbEst, it is computationally expensive, which limits its applicability to large networks. In practice, Myopic is always the second slowest algorithm, after Greedy (Fish_2019, ).
Figure 5. Illustrations of two runs of Myopic for different initial seeds (red), with new selected seeds (yellow), and fixed $\alpha=0.5$. Numbers indicate $\pi$ for each node after the new seed is selected. Initialization significantly affects the performance of Myopic.
@@
Figure 6. Mean performance of the intervention algorithms on each domain in the corpus under medium spreadability. Each algorithm’s performance is averaged over a given domain, 20 runs per network.
Myopic, Naive Myopic, and Gonzalez all start with an initial seed. Past work chose this initial seed to be the highest degree node. Our initial experiments indicate that this choice confers a substantial advantage to these methods (Figure 5), and that some of the previous positive results are thus attributable to this initial seed choice rather than to the algorithms’ subsequent choices. To mitigate this bias we instead initialize all heuristics with a seed set composed of a single uniformly randomly selected initial seed. Moreover, for each evaluation round on a particular network, we initialize all algorithms to use the same random seed, which further controls for the effects of different initial seed choices. This initial seed choice is not counted against the budget $k$ (Figure 4).
3.2. New Fast Algorithms
In addition to the four heuristics from prior work, here we introduce 10 new heuristics to maximize the minimum $\pi_{i}$ on a budget. These heuristics are designed to be computationally lightweight, scaling to far larger networks, while also matching or exceeding the performance of Myopic on the corpus. We group the new algorithms into three families: BFS-based, PPR-based and Topology-based.
@@ 3.2. New Fast Algorithms
While this approach does not exactly measure $\pi_{i}$, it is much faster than ProbEst, taking $\mathcal{O}(n\cdot\langle k\rangle)$ per iteration, because for most networks, including those in our corpus, the mean degree $\langle k\rangle\ll 1000$ (Figure 2). The key design principle of this algorithm is to capture complex network structures that Gonzalez could not account for. In Figure 4, we find that Myopic BFS almost matches Myopic’s performance in situations where Gonzalez lags behind.
@@
In the PPR-based heuristics, Myopic PPR and Naive Myopic PPR, we use Personalized PageRank (PPR) to estimate $\pi_{i}$ instead of ProbEst or BFS. Personalized PageRank (implemented with networkx (SciPyProceedings_11, )) performs a random walk that probabilistically restarts from nodes in the seed set (PPR1998, ). The ranking produced by PPR is not a direct estimate of $\pi_{i}$; we treat the PPR values as being correlated, such that a lower PPR score is a proxy for a lower $\pi_{i}$ value.
Topology-based: LeastCentral and MinDegree.
@@
These heuristics are based on the intuition that nodes with low $\pi_{i}$ values have distinctive structural positions or patterns of connectivity. LeastCentral selects the non-seed node $i$ with the lowest closeness centrality $c_{i}$ as the next seed. Similarly, LeastCentral_n selects the lowest centrality node $i$’s highest degree neighbor as the next seed. Here, the closeness centrality of a node $i$ is the inverse of the average shortest path length from $i$ to all other reachable nodes in the same connected component (FREEMAN1978215, ; newman18, ). Lower closeness centrality implies the node is less reachable by the rest of the network, and therefore is expected to have lower $\pi_{i}$.
@@
The four remaining topology-based heuristics exploit a network’s degree structure to make decisions, based on the observation that in practice Myopic tends to select nodes with low degree as seeds. The first two heuristics take a non-seed node with the lowest degree, breaking ties by choosing the node with the lowest harmonic centrality (boldi2013axioms, ; newman18, ) among same-degree nodes. In the case of MinDegree_hc, it chooses that node itself as the next seed, while MinDegree_hcn chooses the highest-degree neighbor of that node. The two remaining variations replace harmonic centrality in the aforementioned logic with the neighbor degree, i.e. breaking ties by choosing the node with the highest neighbor degree (sum of degrees across neighbors). These heuristics are called MinDegree_nd and MinDegree_ndn, respectively.
4. Experimental Results
@@
Figure 7. Performance of intervention algorithms on the network corpus, relative to Myopic and sorted in ascending order by network size within each domain. Under medium spreadability, $24\%$ of networks have no algorithm better than or within $80\%$ of Myopic’s performance. “Equivalent” is defined as within one standard error of $\beta$ for Myopic; typically about 0.001.
4.1. Performance of Algorithms on the Corpus
@@
We evaluate and compare the performance of 14 algorithms in total (10 new algorithms and 4 from prior work), applied to all $174$ networks in the corpus. Our code and data are available at https://github.com/TheoryInPractice/FairInfoAccessHeuristics. We are interested in algorithms that operate well on a tight budget and so let $k\in\{1,2,\ldots,10\}$ seed nodes. For each spreadability level (low, medium, high), we produce a $10\times 14\times 174$ matrix, where each entry is the $\beta$ performance of an algorithm after adding $k$ seeds in a network, averaged over $20$ runs. We focus on the medium spreadability regime here, and include results for low and high spreadability in Appendix C. Figure 6 displays the mean performance of each algorithm on networks by domain under medium spreadability (see Figure S1 for low and high spreadability results). Across domains, Myopic produces the best average performance.
@@
To further examine the results, we sort within each domain by network size in ascending order, and then score algorithms as better than, “equivalent” to (within one std. err.), within 80% of, or worse than Myopic (see Figure 7; low and high spreadability results in Figure S2). In this way, we can assess whether the better average performance of Myopic applies to individual networks compared to other algorithms. This experiment reveals that Myopic is not universally the best algorithm for all networks in any spreadability setting; on many individual networks other algorithms perform equivalently or better. In the medium spreadability setting, for only $24\%$ of networks is there no algorithm that performs at least $80\%$ as well as Myopic (Figure 7).
@@
This variability in performance across networks suggests that network structure plays a critical role in governing the relative performance of different algorithms. Figure 8 plots the instances for which each algorithm was the best performing algorithm on a network against that network’s mean degree $\langle k\rangle$. Myopic tends to perform better on networks with lower average degree, although it also performs well on many networks with larger mean degree. In practice, we find that the final seed set produced by Myopic is often composed primarily of low-degree nodes located in a network’s periphery, and these nodes may be too far removed from other disadvantaged nodes to meaningfully improve their $\pi_{i}$ values.
@@
A second takeaway is that a few specific algorithms tend to perform better than Myopic in certain settings (Figure 8), specifically MinDegree_hcn and Gonzalez. The performance of MinDegree_hcn in particular tends to improve over Myopic with increasing average degree, while Gonzalez does best in networks with very low mean degrees. Furthermore, we note that in Figure 1(a), on average MinDegree_hcn outperforms Myopic in the economic domain, and from Table S1, we see that the economic domain has the highest mean degrees. The MinDegree_hcn algorithm selects as seeds the highest-degree neighbors of low-degree nodes. As result, new information cascades seeded at these nodes will tend to spread quickly to a number of disadvantaged nodes in the network.
@@
Figure 8. Best-performing algorithm vs. mean degree of the network (medium spreadability), for all networks. Counts on the right show total circles per line, i.e., the number of times an algorithm was the best over the whole corpus.
4.2. Algorithm Runtime
@@
All of the new algorithms are substantially faster than Myopic and Naive Myopic (Figure 9). Because they only require sorting two lists, MinDegree_ndn and MinDegree_nd are the most efficient, improving running times over ProbEst-based algorithms by a factor of 1000-10000x, depending on the size of the network (Figure 9). BFS-based algorithms are marginally slower than these fastest algorithms, and algorithms that use an All-Pairs-Shortest-Paths (APSP) calculation fall between the ProbEst and BFS algorithm groups. As expected from the asymptotics, BFS algorithms tend to scale more slowly (in terms of runtime) with network size than do ProbEst algorithms, while APSP algorithms scale more quickly (Figure 9).
The low upfront cost of APSP-based algorithms makes them far faster than ProbEst-based algorithms in many practical settings, being 10-100x faster on networks with less than $n=10^{6}$. However, the asymptotic cost of APSP is cubic in the network size, implying that for sufficiently large networks, ProbEst will be faster. For our corpus, we estimate the crossover point when Myopic becomes more efficient than MinDegree_hc to occur between $1,262,000$ and $1,515,000$ nodes (95% CI, from 1000 bootstraps). However, approximation algorithms for APSP could potentially extend their practical efficiency much further.
Figure 9. Runtime of old and new algorithms across the network corpus with $k=10$ seeds averaged over 10 runs, showing a substantial advantage in running time for the new algorithms. Note: algorithms with * require ProbEst.
4.3. A Fast Meta-Learning Algorithm